3 Required Information The following information applies to the questions displayed below. Simon Companys year-end balance sheets follow At December 31 Part 3 of 4 2017 2016 2015 Assets Cash Accounts receivable, net Merchandise inventory Prepaid expenses Plant assets, net Total assets $ 35,323 41,290 41,745 96,800 62,700 57,600 111,500 84,500 53,e00 11,375 10,839 4,638 338,764 306,673 256,417 $586,962 $506,082 $413,400 10 points Liabilities and Equity Accounts payable Long-term notes payable secured by $149,077 S 87,225 s 55,660 Hint mortgages on plant assets Common stock, $10 par value Retained earnings Total liabilities and equity 112,556 118,788 93,189 162,500 162,500 162,580 162,829137,569 182,851 $586,962 $506,002 $413,400 The companys income statements for the years ended December 31, 2017 and 2016, follow. Assume that a credit For Year Ended December 31 2017 2016 Sales Cost of goods sold Other operating expenses Interest expense Income taxes Total costs and expenses Net income $763,051 $602,142 $465,461 236,546 12,972 9,920 $391,392 152,342 13,849 9,032 724,899 $ 38,152 $ 2.35 566,615 S 35,527 $ 2.19 Earnings per share (3) Compute inventory turnover 2017 times 2016 times
Answer #1

Inventory turnover ratio =cost of goods sold/Average inventory

2017 =

Average inventory =(111500+84500)/2=$98000

cost of goods sold=$465461

Inventory turnover ratio =465461/98000=4.75 times

2016

Average inventory =(53000+84500)/2=$68750

cost of goods sold=$391392

Inventory turnover ratio=391392/68750=5.69 times